OpenStudy (anonymous):

WILL MEDAL Solve the equation: -3 2/5 + b = 8 1/5

OpenStudy (trojanpoem):

\[-3\frac{ 2 }{ 5 } + b = 8\frac{ 1 }{ 5 }\] b = 11 3/5

OpenStudy (anonymous):

can you help me with 2 more?

OpenStudy (trojanpoem):

Sure.

OpenStudy (anonymous):

-7/3n= - 42

OpenStudy (anonymous):

n=?

OpenStudy (trojanpoem):

n = -42 * -3/7

OpenStudy (trojanpoem):

n = 18 ( if -7 over 3 n)

OpenStudy (anonymous):

oh. ok

OpenStudy (anonymous):

Carla's plant was 10 2/3 inches tall. Yolanda's was 8 3/4 inches tall. How much taller was Carla's plant than Yolanda's?

OpenStudy (trojanpoem):

Get the difference between the two numbers

OpenStudy (anonymous):



OpenStudy (anonymous):

11/12 or 5/12 or 1/12

OpenStudy (trojanpoem):

23/12 = 1 11/12

OpenStudy (anonymous):

oh. i understand now. ok thank you sooo much!!!

OpenStudy (trojanpoem):

You're always welcome.